OpenStudy (anonymous):

A science class has a ratio of 3 boys to 2 girls. What is the largest class that could be formed with less than 40 students?

OpenStudy (anonymous):

The answer would vary

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

a science class has a ratio of 3 boys to 2 girls. B/G = 3/2 | Cross Multiplying 3G = 2B G = (2/3)B Question states*** what is the largest class that could be formed with less than 40 students? B + (2/3)B = 40 for ex B = (3/5)*40 18 and 12 B = (3/5)35 21 and 14 35 the largest class to maintain the ratio